OpenStudy (anonymous):

A Recipe for a dozen cookies calls for 1/4 cup of milk, 2 2/3 cups of flour, and 3 1/2 cups of other ingredients. How many cups of milk,flour and other ingredients are needed to make 48 cookies?

OpenStudy (anonymous):

To get 48 cookies you need to multiply all of the ingredients by 4: (1/4) * 4 = 1 cup of milk (2+(2/3)) * 4 = 10 + 2/3 cups of flower (3+(1/2)) * 4 = 14 cups of other ingredients